OpenStudy (anonymous):

Maria's age is 15 years less than Orlando's age The sum of their age is 58. What are their ages?

OpenStudy (dumbcow):

M=O-15 M+O=58 use substitution (O-15)+O = 58 2O -15 = 58 2O = 73 O = 36.5 solve for M M = 36.5-15 M = 21.5